OpenStudy (anonymous):

e this formula to solve this problem: Power (P) = Work (W) ÷ time (t) Joachim's body supplies a force of 450 N to run up a 20 meter hill in 15 seconds. How much power is involved in Joachim's run up the hill? A. 30 Watts B. 22.5 Watts C. 300 Watts D. 600 Watts

OpenStudy (anonymous):

*use this formula

OpenStudy (thadyoung):

Use it. It's given to you. Just plug and solve.

OpenStudy (thadyoung):

What seems to be the problem here, it's very straight forward.

OpenStudy (anonymous):

P= w/t work(W)= mgh= force*distance W= 450*20=9000 so, P= W/t = 9000/15 = 600 watt